OpenStudy (anonymous):

What's the missing number? 5, 11, 29, 83, ___

OpenStudy (anonymous):

Um, help? :/

OpenStudy (turingtest):

They look prime, this is weird to me.

OpenStudy (mertsj):

They are prime

OpenStudy (anonymous):

That's what I thought..

OpenStudy (anonymous):

245 11-5 = 6 29-11 = 18 83-29 = 54 6 18 54 are geometric series so the next is 54 * 3 = 162 missing number is 83 + 162 = 245

OpenStudy (mertsj):

I think the pattern is triple the number and subtract 4 to get the next number.

OpenStudy (anonymous):

Thanks @adham and @mertsj! :)

OpenStudy (anonymous):

:)

OpenStudy (mertsj):

yw